# Real-World Evaluation of two Cooperative Intersection Management Approaches

### v5

@@ Abstract
Index Terms:
Connected automated driving, cooperative planning, driver model, scenario prediction, optimization, reinforcement learning, graph neural network, mixed traffic, simulation, real-world evaluation, traffic efficiency.white{textblock*}7.1in(0.7in,0.15in)
SUBMITTED TO REVIEW AND POSSIBLE PUBLICATION. COPYRIGHT WILL BE TRANSFERRED WITHOUT NOTICE.
Personal use of this material is permitted.
Permission must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
Urban traffic is prone to inefficiencies and disturbances due to the ever-increasing volume of traffic. This manifests, e.g., at smaller intersections, where static priority rules are the prevailing method for coordination of vehicles. Connected automated driving opens up new opportunities to improve urban traffic efficiency by leveraging communication links between vehicles and possibly infrastructure systems. Moreover, edge computing resources are becoming available in urban areas that enable, e.g., populating and maintaining of a collective environment model (EM) of an area of interest. This setup allows connected automated vehicles (CAVs) to improve their local planning algorithms by incorporating data from the server-side EM, as it was shown in [4] at a suburban three-way intersection in Ulm-Lehr, Germany.
@@
Our methods are designed for urban intersections with one lane per direction. Larger intersections are beyond the scope of this paper, as they are typically managed by traffic lights and require a different cooperative planning paradigm.
@@ II-A Classical Cooperative Maneuver Planning
Fewer researchers have tackled the intersection management task in the much more challenging mixed traffic setting and especially scenarios with a vast majority of HDVs, which will be prevalent in the near future. To the best of our knowledge, all previous mixed traffic approaches need to incorporate traffic lights to support the presence of HDVs. The first notable approaches [7, 2] were designed for a low rate of HDVs ( 10 %) and the gains deteriorate in an environment with a higher amount of HDVs. Two later publications [29, 43] proposed virtual platoons for signalized intersection management, supporting mixed traffic at high HDV rates and being evaluated in a real-time traffic simulation. However, all methods mentioned above require traffic lights to control HDVs and as such are out of our problem scope. Also, most approaches assume that the turn direction of HDVs is known, which is generally not given in real-world traffic.
@@ III-A Definition and Coordination of Cooperative Maneuvers
For both planning modules, the maneuvers are passed to the CAVs according to the maneuver coordination protocol proposal in [26]. Depending on the number of vehicles on conflicting paths, up to two maneuver waypoints (one at the intersection entry and one at the intersection exit) are passed to a given CAV. Each vehicle remains responsible for its own motion planning but needs to fulfill constraints defined by the maneuver waypoint as
The interval denotes the admitted time window for the vehicle to cross the waypoint . The fields and describe the preceding and following vehicle IDs, respectively. This additional information can be leveraged by vehicle-side motion planners to improve follow trajectory planning on lead vehicles [32]. Note that lateral guidance is provided by lane centerlines in a common map that is available to all vehicle-side motion planners.
@@ III-B Environment Model Processing
The basis for both cooperative maneuver planning modules is the server-side EM. This EM is the result of fusing multiple data sources, like cooperative awareness messages (CAM, [8]) and collective perception messages (CPM, [9]) from connected vehicles or infrastructure perception [4] and is considered to contain all information relevant for maneuver planning. A vehicle in the EM is denoted as
where denotes a unique identifier and the binary flag determines whether the vehicle is a CAV and thus controllable. describes the vehicle’s pose on the local 2D ground plane and is the current driving speed. CAVs share their intended route or destination , which is unknown for HDVs, though, which are assumed to be non-connected.
@@ III-C2 Priority Assignment Search
Each possible maneuver is internally represented as a list of priority assignments between pairs of cooperative vehicles. Each entry is a pair of vehicles , where the first is prioritized over the second. The planner operates cyclically in a frequency of 5 Hz and tries to add assignments to the previous cycle result to optimize the efficiency, by predicting the scenario for each .
@@ III-D1 Learning Model
Due to the lack of ground-truth data for cooperative maneuvers in urban traffic, we consider the planning task a multi-agent RL problem. The joint planning task is modeled by a single partially observable Markov decision process. Thus, the dimensionality of the observation space depends on the number of vehicles currently in the scene and may vary over time. Moreover, the dimensionality of the action space depends on the number of cooperative vehicles present and requires a permutation equivariant mapping to the observation space.
@@ III-D2 Graph-based Representation and Network
The graph based scene representation used in this study is based on our proposal in [19], which proved to be suited for learning a sensible mixed traffic capable RL policy. Thus, the current traffic scene observation is defined as a directed graph, as illustrated in Fig. 3. The nodes of vehicles that are in conflict and thus need to be coordinated are connected by directed edges. Due to the inherently unknown turning intention of HDVs, like in the figure, the graph comprises additional edges for all conflicts that cannot be ruled out reliably. Different kinds of interactions are encoded by means of edge types. The crossing edge type indicates a pair of vehicles that are located in front of the intersection driving on paths that intersect or merge on the intersection area. Leader-follower relations are encoded by an edge of type same lane pointing from the leader to the following vehicle.
@@ III-E Prediction and Derivation of Cooperative Maneuvers
For the optimization-based planner, the selected priority assignments need to be converted to maneuver constraints as in Eq. (1) for each individual CAV. Each pair of conflicting CAVs shares a conflict zone, i.e., a segment along the road that may only be occupied by one of the vehicles at any time to guarantee freedom from collision, as explained above. The predicted timestamp at which the prioritized vehicle leaves and the yielding vehicle may enter the conflict zone is used to encode the space-time maneuver constraints. The start and end points of the conflict zone along the respective vehicle routes are used as waypoints for the constraints.
The RL-based cooperative planner derives cooperative maneuvers using a built-in simulator. After initializing the simulator based on the current state of the server-side EM, the RL policy is queried to predict the future evolution of the scene. For each recorded simulated trajectory, waypoint candidates are dynamically created at the entry and exit of the intersection conflict zone. Similar to the optimization-based planner, the lower and upper time interval bounds are obtained from the simulation. Thereby, the entry waypoint determines the earliest possible time of entering the intersection, while the waypoint at the intersection exit specifies when the receiving CAV shall clear the conflict area at the latest.
@@ IV-A Simulation Setup
The simulation framework that we used for evaluation of the planning methods builds upon the DeepSIL simulator introduced in [36]. It is based on the ROS2 software distribution [24] and can simulate HDVs as well as CAVs. We extended the framework to support multiple CAVs and cooperative maneuvers. The simulation has a base time step of 50 ms and runs in wall-clock time on a single machine. This means that processing delays from the cooperative maneuver planners are considered and affect the results, while communication delays are not modeled. Due to limited computational resources on the machine, simulations with more than four CAVs are slowed down slightly to still be able to run all vehicle trajectory planners in parallel.
@@
Connected automated vehicles are simulated using a single-track model with friction and dead time, resembling one of our real-world CAVs. The inputs to this model are computed by the trajectory planner and controller from [32], which is part of this CAV’s software and therefore provides a realistic CAV behavior. Like HDVs, CAVs do not know the route of other vehicles, even of other CAVs, and assume the most conflicting route. However, the simulated CAVs do not need to approach the intersection slowly due to occlusion, because they receive the intersection EM as in [4].
@@ IV-B Scenario Definition
We define all simulated scenarios at an exemplary intersection in Ulm-Lehr, Germany, as shown in Fig. 4, where also our pilot site (cf. Section VI-A) is located. It is a T-junction with a bending main road and a minor subordinate road. The speed limit is () on the eastern and western part and () on the northern part.
Within this setting, we created a total set of 480 scenarios that serve as an initial configuration to the simulation. In each scenario, a number of vehicles is spawned at the locations shown in Fig. 4, where each vehicle has an initial speed of and a specific route, i.e., turn direction. The first 280 scenarios are an enumeration of all combinations of source and target directions with up to two vehicles from each direction and in which there is at least one pair of conflicting vehicles. The remaining 200 scenarios are randomly sampled with up to four vehicles from each direction and a random route, limited to eight vehicles total due to computational constraints.
The scenario set was simulated both in fully automated traffic and mixed traffic. When simulating mixed traffic, the vehicles are randomly divided into 50 % HDVs and 50 % CAVs, where for an odd number of vehicles, the last one is chosen randomly. Out of the 480 total scenarios, less than 3 % of the simulations were not successful and resulted in a timeout or software crash (five scenarios in fully automated traffic, 14 scenarios in mixed traffic). This is due to remaining subtle issues in the prototype CAV software stack and unrelated to the maneuver planning. Those scenarios were excluded from the evaluation in all maneuver planning configurations.
@@ V-A Effectiveness and Efficiency Analysis
Fig. 5 depicts the results in fully automated traffic. Both planners are able to realize cooperative maneuvers in the vast majority of scenarios, as can be seen from Fig. LABEL:sub@sfig:all_cav_deviating. For comparability, the analyses in this section are performed over all cooperative scenarios, that is, all scenarios in which at least one cooperative planner deviates from the baseline crossing order. For all other scenarios, the metrics do not significantly differ from the baseline. In the cooperative scenarios, the traffic flow especially on the subordinate road is improved significantly, while the optimization-based planner provides an additional benefit over the RL-based one. Fig. LABEL:sub@sfig:all_cav_relative_duration_road_type depicts the distribution of the duration required by the vehicles to complete the maneuver relative to non-cooperative CAVs. Per definition, NC traffic exhibits a relative duration of and is thus not shown in the plot. Almost all vehicles on the subordinate road benefit of a reduced delay due to interaction. Some vehicles on the major road take slightly longer to complete the maneuver in the cooperative case. This is expected since the vehicles on the major road waive their unconditional right of way when taking part in cooperative maneuvers. Overall, the cooperation leads to a reduction in median duration, which becomes even more pronounced when comparing to pure HDV traffic. The optimization-based planner achieves slightly higher efficiency due to planning with a lower safety distance, as discussed in Section V-B.
@@
In mixed traffic, both cooperative planning approaches increase the traffic efficiency, as illustrated by Fig. 6, although to a lesser extent than in fully automated traffic. It is worth noting that a 50 % share of HDVs renders many cooperative maneuvers infeasible because the priority relations towards HDVs have to be retained. Thus, achieving a notable gain in efficiency is much more challenging in mixed traffic, which is consistent with the individual planner evaluations [25, 19].
@@ V-B Criticality Analysis
The criticality results for fully automated traffic and for mixed traffic are depicted in Table II and Table III, respectively. Both planner methods yield a higher share of critical scenarios compared to the baseline, which is expected as they are optimized for efficiency. The RL planner results are closer to the NC baseline, which is most likely due to the strong focus on HDV interoperability in the reward scheme and training procedure. Meanwhile, the optimization-based planner shows significantly increased criticality percentages for DRAC and PET, which goes along with its better performance in traffic efficiency metrics. This is due to the fact that, internally, the planner calculates geometrically accurate conflict zones and thus assumes a PET of 0 s to be sufficiently safe. Increasing this safety distance to allow for more reaction time will be considered in future works.
@@ VI-A Pilot Site Ulm-Lehr
The real-world experiments were conducted at a pilot site for connected automated driving in Ulm-Lehr, Germany [4]. An aerial view of the site is shown in Fig. 7. The unsignalized intersection is equipped with multiple sensor processing units, each consisting of cameras, radar and lidar sensors as well as a computer for object detection. A fusion and tracking module [13] on an edge server combines all object detections into the infrastructure EM, which is provided to the connected vehicles and the cooperative planning modules on the same edge server. We employed three CAVs, one of which used the trajectory planning from [32] while the other two implemented the concept from [38]. The CAVs, infrastructure, and edge server are connected via a 5G SA cellular network with a communication latency of about 15 ms.

### v6

@@ I Introduction
Urban traffic is prone to inefficiencies and disturbances due to the ever-increasing volume of traffic [16]. This manifests, e.g., at smaller intersections, where static priority rules are the prevailing method for coordination of vehicles. Connected automated driving opens up new opportunities to improve urban traffic efficiency by leveraging communication links between vehicles and possibly infrastructure systems. Moreover, edge computing resources are becoming available in urban areas that enable, e.g., populating and maintaining of a collective environment model (EM) of an area of interest. This setup allows connected automated vehicles (CAVs) to improve their local planning algorithms by incorporating data from the server-side EM, as it was shown in [4] at a suburban three-way intersection in Ulm-Lehr, Germany.
@@
Our methods are designed for urban intersections with one lane per crossing direction. Larger intersections are beyond the scope of this paper, as they are typically managed by traffic lights and require a different cooperative planning paradigm.
@@ II-A Classical Cooperative Maneuver Planning
Fewer researchers have tackled the intersection management task in the much more challenging mixed traffic setting and especially scenarios with a vast majority of HDVs, which will be prevalent in the near future. To the best of our knowledge, all previous mixed traffic approaches need to incorporate traffic lights to support the presence of HDVs. The first notable approaches [7, 2] were designed for a low rate of HDVs ( 10 %) and the gains deteriorate in an environment with a higher amount of HDVs. Two later publications [30, 44] proposed virtual platoons for signalized intersection management, supporting mixed traffic at high HDV rates and being evaluated in a real-time traffic simulation. However, all methods mentioned above require traffic lights to control HDVs and as such are out of our problem scope. Also, most approaches assume that the turn direction of HDVs is known, which is generally not given in real-world traffic. Prior works claiming HDV support without traffic lights or other means of communication with human drivers depend on strong assumptions on the HDVs’ behavior in mixed traffic. The reservation-based approach [45] assigns CAVs a priority based on their lane, while expecting HDVs to follow a possibly present leading CAV closely. The application of such an approach in public traffic where HDVs are not specifically instructed remains questionable.
@@ III-A Definition and Coordination of Cooperative Maneuvers
For both planning modules, the maneuvers are passed to the CAVs according to the maneuver coordination protocol proposal in [27]. Depending on the number of vehicles on conflicting paths, up to two maneuver waypoints (one at the intersection entry and one at the intersection exit) are passed to a given CAV. The CAVs need to fulfill constraints defined by the maneuver waypoint as
The interval denotes the admitted time window for the vehicle to cross the waypoint . The fields and describe the preceding and following vehicle IDs, respectively. This additional information can be leveraged by vehicle-side motion planners to improve follow trajectory planning on lead vehicles [33]. Each CAV remains responsible for its own motion planning and validation whether the maneuver constraints can be safely incorporated. Note that lateral guidance is provided by lane centerlines in a common map that is available to all vehicle-side motion planners.
@@ III-B Environment Model Processing
The basis for both cooperative maneuver planning modules is the server-side EM. This EM is the result of fusing multiple data sources, like cooperative awareness messages (CAM, [8]) and collective perception messages (CPM, [9]) from connected vehicles or infrastructure perception [4]. It is considered to contain all information relevant for maneuver planning with full observation of the intersection area. A vehicle in the EM is denoted as
where denotes a unique identifier and the binary flag determines whether the vehicle is a CAV and thus controllable. describes the vehicle’s pose on the local 2D ground plane and is the current driving speed. CAVs share their intended route or destination as a set of sparse waypoints, which is unknown for HDVs, though, which are assumed to be non-connected.
@@ III-C2 Priority Assignment Search
Each possible maneuver is internally represented as a list of priority assignments between pairs of cooperative vehicles. Each entry is a pair of vehicles , where the first is prioritized over the second. The planner operates cyclically in a frequency of 5 Hz and tries to add assignments to the previous cycle result to optimize the efficiency, by predicting the scenario for each . As the future behavior of HDVs, such as gap acceptance and turning direction, is unknown and cannot be controlled, all probable actions of HDVs have to be considered. Thus, a dependency graph analysis only considers maneuvers independent of HDV behavior as valid, largely reducing the number of feasible cooperative maneuvers in mixed traffic.
@@
The number of possible priority assignments and, therefore, the simulation and search run time rises with the number of conflicting CAV pairs. To remain real-time capable in larger scenarios, the search is limited to the first 100 priority assignments, keeping the planning time well below . This limit was, however, not reached in the considered scenarios.
@@ III-D1 Learning Model
Due to the lack of ground-truth data for cooperative maneuvers in urban traffic, we consider the planning task a multi-agent RL problem. The joint planning task is modeled by a single partially observable Markov decision process, since the driving intention and destination of the HDVs is not known to the planning algorithm. Thus, the dimensionality of the observation space depends on the number of vehicles currently in the scene and may vary over time. Moreover, the dimensionality of the action space depends on the number of cooperative vehicles present and requires a permutation equivariant mapping to the observation space.
@@ III-D2 Graph-based Representation and Network
The graph based scene representation used in this study is based on our proposal in [20], which proved to be suited for learning a sensible mixed traffic capable RL policy. Thus, the current traffic scene observation is defined as a directed graph, as illustrated in Fig. 3. The nodes of vehicles that are in conflict and thus need to be coordinated are connected by directed edges, while it is assumed that vehicles follow and stay within their lanes. Due to the inherently unknown future behavior of HDVs, including gap acceptance and turning intention, the graph comprises additional edges for all potential conflicts, like with in the figure. Only once the maneuver of the HDV becomes clear (typically while entering the intersection), the superfluous edges are removed and the cooperative maneuver is subsequently optimized. Different kinds of interactions are encoded by means of edge types. The crossing edge type indicates a pair of vehicles that are located in front of the intersection driving on paths that intersect or merge on the intersection area. Leader-follower relations are encoded by an edge of type same lane pointing from the leader to the following vehicle.
@@
Apart from a suited network architecture, RL requires a well-designed reward function to learn a reasonable policy. Details on the definition and tuning of the various reward components can be found in [19, 20].
RL planning runs are triggered regularly With a period of , which facilitates the observation of the CAVs’ reactions to the cooperative maneuver with the start of the next planning run. For typical traffic scenarios at smaller intersections, the planning run takes significantly less than the period length of . In unfavorable circumstances with many (non-connected) vehicles, a full planning run might take longer, which is caught by a timeout that ensures real-time capability [21].
@@ III-E Prediction and Derivation of Cooperative Maneuvers
For the optimization-based planner, the selected priority assignments need to be converted to maneuver constraints as in Eq. (1) for each individual CAV. Each pair of conflicting CAVs shares a conflict zone, i.e., a segment along the road that may only be occupied by one of the vehicles at any time to guarantee freedom from collision, as explained above. The predicted timestamp at which the prioritized vehicle is predicted to leave the conflict zone is used to encode the space-time maneuver constraints for entering and exiting. The start and end points of the conflict zone along the respective vehicle routes are used as waypoints for the constraints.
The RL-based cooperative planner derives cooperative maneuvers using a built-in simulator. After initializing the simulator based on the current state of the server-side EM, the RL policy is queried to predict the future evolution of the scene. Analogously to the optimization-based planner, the crossing points of the simulated trajectory and the intersection conflict zone are used to derive so-called waypoint candidates. Only if there is a conflicting vehicle, the corresponding candidate waypoint is used as a maneuver waypoint. Thus, only if there is a preceding vehicle, the waypoint at the intersection entry is included in the maneuver. Otherwise, the CAV is free to enter the intersection as soon as it considers it beneficial to the maneuver, which leaves room for optimization to the CAV motion planning.
As noted above, the CAVs remain responsible to check the maneuver for local feasibility, and the maneuver planners cannot guarantee that the CAVs can execute the suggested maneuver. However, while the high-level maneuver planning is optimized for speed, the driver models internally used in each maneuver planner strive to resemble realistic behavior, which usually is far from the physical (or even comfortable) limits at intersections. Thus, the coordinated CAVs can usually easily comply with the required time constraints.
@@ IV-A Simulation Setup
The simulation framework that we used for evaluation of the planning methods builds upon the DeepSIL simulator introduced in [37]. It is based on the ROS2 software distribution [25] and can simulate HDVs as well as CAVs. We extended the framework to support multiple CAVs and cooperative maneuvers. The simulation has a base time step of 50 ms and runs in wall-clock time. This means that processing delays from the cooperative maneuver planners are considered and affect the results, while communication delays are not modeled.
We use a single potent simulation machine with an AMD Ryzen 3990X CPU. Still, running multiple instances of large parts of the CAV software stack as well as multiple deep-learning-based HDV models requires a lot of computational resources, while the maneuver planning requirements are comparatively modest. Due to those limitations, the simulations need to be limited to eight vehicles total. Additionally, scenarios with more than four CAVs are slowed down slightly to cope with immaturity in the ROS2 communication layer.
@@
CAVs are simulated using a single-track model with friction and dead time, resembling one of our real-world CAVs. The inputs to this model are computed by the trajectory planner and controller from [33], which is part of this CAV’s software and therefore provides a realistic CAV behavior. Like HDVs, CAVs do not know the route of other vehicles, even of other CAVs, and assume the most conflicting route. However, the simulated CAVs do not need to approach the intersection slowly due to occlusion, because they receive the intersection EM as in [4].
@@ IV-B Scenario Definition
We define all simulated scenarios at an exemplary intersection in Ulm-Lehr, Germany, as shown in Fig. 4, where also our pilot site (cf. Section VI-A) is located for coherent evaluations. It is a T-junction with a bending main road and a minor subordinate road. The speed limit is () on the eastern and western part and () on the northern part.
Within this setting, we created a total set of 480 scenarios that serve as an initial configuration to the simulation. In each scenario, a number of vehicles is spawned at the locations shown in Fig. 4, where each vehicle has an initial speed of and a specific route, i.e., turn direction. The first 280 scenarios are an enumeration of all combinations of source and target directions with up to two vehicles from each direction and in which there is at least one pair of conflicting vehicles. The remaining 200 scenarios are randomly sampled with up to four vehicles from each direction and a random route, limited to eight vehicles total as explained above.
The scenario set was simulated both in fully automated traffic and mixed traffic. When simulating mixed traffic, the vehicles are randomly divided into 50 % HDVs and 50 % CAVs, where for an odd number of vehicles, the last one is chosen randomly. Out of the 480 total scenarios, less than 3 % of the simulations were not successful and resulted in a timeout or software crash (five scenarios in fully automated traffic, 14 scenarios in mixed traffic). Those issues are unrelated to the maneuver planning and were caused by other CAV software modules and immaturity in the ROS2 communication layer. Thus, those scenarios were excluded from the evaluation in all maneuver planning configurations.
@@ V-A Effectiveness and Efficiency Analysis
Fig. 5 depicts the results in fully automated traffic. Both planners are able to realize cooperative maneuvers in the vast majority of scenarios, as can be seen from Fig. LABEL:sub@sfig:all_cav_deviating. The “original” order is the one resulting from simulating without cooperation. In general, the crossing order is not derivable from static priority rules, as gaps in main road traffic can be used. In our specific scenario, however, there are no gaps, so the non-cooperative crossing order is unambiguous and any deviation results from an applied cooperative maneuver.
For comparability, the analyses in this section are performed over all explicitly cooperative scenarios, that is, all scenarios in which at least one cooperative planner deviates from the baseline crossing order. For all other scenarios, the metrics do not significantly differ from the baseline. In the cooperative scenarios, the traffic flow especially on the subordinate road is improved significantly, while the optimization-based planner provides an additional benefit over the RL-based one. Fig. LABEL:sub@sfig:all_cav_relative_duration_road_type depicts the distribution of the duration required by the vehicles to complete the maneuver relative to non-cooperative CAVs. Per definition, NC traffic exhibits a relative duration of and is thus not shown in the plot. Almost all vehicles on the subordinate road benefit of a reduced delay due to interaction. Some vehicles on the major road take slightly longer to complete the maneuver in the cooperative case. This is expected since the vehicles on the major road waive their unconditional right of way when taking part in cooperative maneuvers. Overall, the cooperation leads to a reduction in median duration, which becomes even more pronounced when comparing to pure HDV traffic. The optimization-based planner achieves slightly higher efficiency due to planning with a lower safety distance, as discussed in Section V-B.
@@
In mixed traffic, both cooperative planning approaches increase the traffic efficiency, as illustrated by Fig. 6, although to a lesser extent than in fully automated traffic. It is worth noting that a 50 % share of HDVs renders many cooperative maneuvers infeasible because the priority relations towards HDVs have to be retained. Thus, achieving a notable gain in efficiency is much more challenging in mixed traffic, which is consistent with the individual planner evaluations in [26, 20]. We refer to these publications for more in-depth simulative and dataset-based evaluations with multi-lane intersections.
@@ V-B Criticality Analysis
The criticality results for fully automated traffic and for mixed traffic are depicted in Table II and Table III, respectively. Please note that we evaluated all and not only the deviating scenarios as the analysis of the criticality metrics shall only convey qualitative trends. For this reason, there is no claim about the absolute numbers or in-depth discussion.
Both planner methods yield a higher share of critical scenarios compared to the baseline, which is expected as they are optimized for efficiency. The RL planner results are closer to the NC baseline, which is most likely due to the strong focus on HDV interoperability in the reward scheme and training procedure. Meanwhile, the optimization-based planner shows significantly increased criticality percentages for DRAC and PET, which goes along with its better performance in traffic efficiency metrics. This is due to the fact that, internally, the planner calculates geometrically accurate conflict zones and thus assumes a PET of 0 s to be sufficiently safe. Increasing this safety distance to allow for more reaction time will be considered in future works.
@@ VI-A Pilot Site Ulm-Lehr
The real-world experiments were conducted at a pilot site for connected automated driving in Ulm-Lehr, Germany [4]. An aerial view of the site is shown in Fig. 7. The unsignalized intersection is equipped with multiple sensor processing units, each consisting of cameras, radar and lidar sensors as well as a computer for object detection. A fusion and tracking module [13] on an edge server combines all object detections into the infrastructure EM, which is provided to the connected vehicles and the cooperative planning modules on the same edge server. The server uses an AMD Ryzen 9 3950X CPU, similar to the simulation machine. We employed three CAVs, one of which used the trajectory planning from [33] while the other two implemented the concept from [39]. The CAVs, infrastructure, and edge server are connected via a 5G SA cellular network with a communication latency of about 15 ms.